What is the depreciation deduction, using the declining balance (DB) method with a 200% DB ratio, associated with the third year for an asset that cost $35,000 and has an estimated MV of $9,000 at the end of its seven-year useful life? Express your answer in terms of dollars and cents, rounded to the nearest cent (e.g., 1234.56).
Answer #1

Purchase price = 35000

Salvage value = 9000

useful life = 7 years

200% DB, so d= 2/useful life = 2/7

Depreciation in any year m = Purchase price *(1-d)m-1*d

Depreciation in year 3 = 35000 * (1-2/7)3-1*(2/7)

= 35000*0.510204*2/7

=5102.04